OpenStudy (anonymous):

sqrt (8x + 8) + sqrt (2x + 2)

OpenStudy (anonymous):

i know the answer is 3 sqrt (2x + 2) cuz its in the bak of my book but idk how to get that.

OpenStudy (anonymous):

you may want to try factoring out 2 or something like that out of both square roots

OpenStudy (anonymous):

you may want to check the back of the book again cuz when i did it i got 5*sqrt(2x+2)

OpenStudy (anonymous):

\[\sqrt{8x+8}+\sqrt{2x+2}=\sqrt{4(x+2)}+\sqrt{2x+2}\]\[=\sqrt{4}\sqrt{2x+2}+\sqrt{2x+2}=2\sqrt{2x+2}+\sqrt{2x+2}\]\[=3\sqrt{2x+2}\]